# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ Appendix D $-\ell$ Achievability Analysis for Bernoulli Designs
In this part, we identify the more explicit condition for (C1) under the given $\ell$ and a specific feasible $(C,\zeta)$ with $C\leq C_{0}$. We seek to bound $k_{\ell,C,\zeta}$. Our strategy is to first calculate $\mathbbm{E}k_{\ell,C,\zeta}$ and then deduce the high-probability behaviour of $k_{\ell,C,\zeta}$ via Markov’s inequality. Since $k_{\ell,C,\zeta}=\sum_{\mathcal{J}\subset s,|\mathcal{J}|=\ell}\mathbbm{1}(M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k},\leavevmode\nobreak\ M_{\mathcal{J}0}=\zeta M_{\mathcal{J}})$, we have
$\mathbbm{E}k_{\ell,C,\zeta}$ $=\binom{k}{\ell}\mathbb{P}\left(\text{for fixed }\mathcal{J}\subset s\text{ with }|\mathcal{J}|=\ell,\leavevmode\nobreak\ M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k},\leavevmode\nobreak\ M_{\mathcal{J}0}=\zeta M_{\mathcal{J}}\right)$ (92)
$=\binom{k}{\ell}\mathbb{P}\left(\text{Bin}\big{(}n,P_{1}\big{)}=\frac{Cn\nu e^{-\nu}\ell}{k}\right)\mathbb{P}\left(\text{Bin}\Big{(}\frac{Cn\nu e^{-\nu}\ell}{k},\rho\Big{)}=\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}\right)$ (93)
$\leq\binom{k}{\ell}\exp\left(-n\cdot D\Big{(}\frac{C\nu e^{-\nu}\ell}{k}\big{\|}\frac{\nu e^{-\nu}\ell(1+o(1))}{k}\Big{)}-\frac{Cn\nu e^{-\nu}\ell}{k}D(\zeta\|\rho)\right)$ (94)
$=\binom{k}{\ell}\exp\Big{(}-\frac{n\nu e^{-\nu}\ell}{k}\Big{[}f_{1}(C,\zeta,\rho)+o(1)\Big{]}\Big{)},$ (95)
where in (94) we use the Chernoff bound (see (23) in Lemma 1), and in (95) we define
$f_{1}(C,\zeta,\rho):=C\log C-C+C\cdot D(\zeta\|\rho)+1,$ (96)
and then use $D(\frac{C\nu e^{-\nu}\ell}{k}\|\frac{\nu e^{-\nu}\ell(1+o(1))}{k})=\frac{\nu e^{-\nu}\ell}{k}(C\log C-C+1+o(1))$. We note that $f_{1}(C,\zeta,\rho)$ coincides with (56) in the proof of the converse bound for Bernoulli designs.
@@
$\mathbb{P}\Big{(}k_{\ell,C,\zeta}\geq(\ell\log k)^{5}\mathbbm{E}[k_{\ell,C,\zeta}]\Big{)}\leq(\ell\log k)^{-5}.$ (97)
Using the bound on $\mathbbm{E}k_{\ell,C,\zeta}$ from (95), we know that with probability at least $1-(\ell\log k)^{-5}$, it holds that
$k_{\ell,C,\zeta}<(\ell\log k)^{5}\mathbbm{E}k_{\ell,C,\zeta}\leq$ $(\ell\log k)^{5}\binom{k}{\ell}\exp\Big{(}-\frac{n\nu e^{-\nu}\ell}{k}\Big{[}f_{1}(C,\zeta,\rho)+o(1)\Big{]}\Big{)}$ (98)
$\leq$ $\exp\left((1+o(1))\ell\log\frac{k}{\ell}-\frac{n\nu e^{-\nu}\ell}{k}\Big{[}f_{1}(C,\zeta,\rho)+o(1)\Big{]}\right).$ (99)
@@
$\mathbb{P}\Big{(}k_{\ell,C,\zeta}\geq k^{5}\mathbbm{E}[k_{\ell,C,\zeta}]\Big{)}\leq k^{-5}.$ (100)
Combining with (95), we know that with probability at least $1-k^{-5}$, it holds that
$k_{\ell,C,\zeta}<k^{5}\mathbbm{E}k_{\ell,C,\zeta}\leq$ $k^{5}\binom{k}{\ell}\exp\Big{(}-\frac{n\nu e^{-\nu}\ell}{k}\Big{[}f_{1}(C,\zeta,\rho)+o(1)\Big{]}\Big{)}$ (101)
$\leq$ $\exp\Big{(}(1+o(1))\ell\log\frac{k}{\ell}-\frac{n\nu e^{-\nu}\ell}{k}\Big{[}f_{1}(C,\zeta,\rho)+o(1)\Big{]}\Big{)}.$ (102)
@@
Combining the above two cases, we obtain the following conclusion for given $(\ell,C,\zeta)$:
$\frac{n\nu e^{-\nu}\ell}{k}\big{[}f_{1}(C,\zeta,\rho)+o(1)\big{]}\geq(1+\eta_{1})\ell\log\frac{k}{\ell}\leavevmode\nobreak\ \text{for some }\eta_{1}>0$ (103)
@@
$\Longrightarrow$ $\leavevmode\nobreak\ \mathbb{P}\big{(}k_{\ell,C,\zeta}\geq 1\big{)}\leq\hat{P}_{\ell},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where }\hat{P}_{\ell}=\begin{cases}(\ell\log k)^{-5},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }1\leq\ell\leq\log k\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ k^{-5},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }\log k<\ell\leq\frac{k}{\log k}\end{cases},$ (104)
@@
where we implicitly assume that $k$ is sufficiently large. Therefore, if (103) holds, then (C1) holds for some given $(\ell,C,\zeta)$ with probability at most $\hat{P}_{\ell}$.
Condition for (C2) under given $(\ell,C,\zeta)$
@@
We again consider a given $\ell$ and a specific feasible pair of $(C,\zeta)$ with $C\leq C_{0}$, and now switch to analyzing (C2). For $\mathcal{J}^{\prime}\subset[p]$ with $|\mathcal{J}^{\prime}|=\ell$, the probability of a test containing some item from $\mathcal{J}^{\prime}$ is:
$P_{2}:=1-\big{(}1-\frac{\nu}{k}\big{)}^{\ell}=\frac{\nu\ell(1+o(1))}{k}.$ (105)
@@
Given $(\ell,C,\zeta)$, (C2) states that (84) holds for some $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and some $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$. In (84), recall that $G_{\mathcal{J},\mathcal{J}^{\prime},1}$ ($G_{\mathcal{J},\mathcal{J}^{\prime},2}$, resp.) is the number of tests in $\mathcal{N}_{01}\cup\mathcal{M}_{\mathcal{J}1}$ ($\mathcal{N}_{00}\cup\mathcal{M}_{\mathcal{J}0}$, resp.) that contain some item from $\mathcal{J}^{\prime}$, and it will be convenient to further decompose these into
$G_{\mathcal{J},\mathcal{J}^{\prime},1}=\widetilde{G}_{\mathcal{J}^{\prime},1}+U_{\mathcal{J},\mathcal{J}^{\prime},1},\leavevmode\nobreak\ G_{\mathcal{J},\mathcal{J}^{\prime},2}=\widetilde{G}_{\mathcal{J}^{\prime},2}+U_{\mathcal{J},\mathcal{J}^{\prime},2},$ (106)
@@
where we define $\widetilde{G}_{\mathcal{J}^{\prime},1}$ ($\widetilde{G}_{\mathcal{J}^{\prime},2}$, resp.) as the number of tests in $\mathcal{N}_{01}$ ($\mathcal{N}_{00}$, resp.) that contain some item from $\mathcal{J}^{\prime}$ (thus $\widetilde{G}_{\mathcal{J}^{\prime},1}$ and $\widetilde{G}_{\mathcal{J}^{\prime},2}$ have no dependence on $\mathcal{J}$, as also reflected by our notation), and $U_{\mathcal{J},\mathcal{J}^{\prime},1}$ ($U_{\mathcal{J},\mathcal{J}^{\prime},2}$, resp.) as the number of tests in $\mathcal{M}_{\mathcal{J}1}$ ($\mathcal{M}_{\mathcal{J}0}$, resp.) that contain some item from $\mathcal{J}^{\prime}$. Given $(\mathbf{X}_{s},\mathbf{Y})$ and for specific $\mathcal{J}$ and $\mathcal{J}^{\prime}$, it follows that
$\widetilde{G}_{\mathcal{J}^{\prime},1}\sim\text{Bin}(N_{01},P_{2}),\leavevmode\nobreak\ \widetilde{G}_{\mathcal{J}^{\prime},2}\sim\text{Bin}(N_{00},P_{2})$ (107)
$U_{\mathcal{J},\mathcal{J}^{\prime},1}\sim\text{Bin}(M_{\mathcal{J}1},P_{2}),\leavevmode\nobreak\ U_{\mathcal{J},\mathcal{J}^{\prime},2}\sim\text{Bin}(M_{\mathcal{J}0},P_{2})$ (108)
@@
are independent, as the tests comprising $\mathcal{N}_{01},\mathcal{N}_{00},\mathcal{M}_{\mathcal{J}0},\mathcal{M}_{\mathcal{J}1}$ are disjoint. Under these definitions (84) can be written as
$\underbrace{\widetilde{G}_{\mathcal{J}^{\prime},1}-\widetilde{G}_{\mathcal{J}^{\prime},2}}_{:=\widetilde{G}_{\mathcal{J}^{\prime}}}+U_{\mathcal{J},\mathcal{J}^{\prime},1}-U_{\mathcal{J},\mathcal{J}^{\prime},2}\geq(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k},$ (109)
@@
$\max_{\begin{subarray}{c}\mathcal{J}^{\prime}\subset[p]\setminus s\\ |\mathcal{J}^{\prime}|=\ell\end{subarray}}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k}-\max_{\begin{subarray}{c}\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}\\ \mathcal{J}^{\prime}\subset[p]\setminus s,|\mathcal{J}^{\prime}|=\ell\end{subarray}}\leavevmode\nobreak\ U_{\mathcal{J},\mathcal{J}^{\prime},1}.$ (110)
Thus, to bound the probability of (C2), it suffices to bound the probability of its necessary condition given in (110).
@@
The effect of $U_{\mathcal{J},\mathcal{J}^{\prime},1}$: For convenience, we write the last term in (110) using the shorthand $\max_{\mathcal{J},\mathcal{J}^{\prime}}U_{\mathcal{J},\mathcal{J}^{\prime},1}$ with the constraints left implicit. We first show that the term $\max_{\mathcal{J},\mathcal{J}^{\prime}}U_{\mathcal{J},\mathcal{J}^{\prime},1}$ in the right-hand side of (110) only has a minimal impact, in the sense that scales as $o(\frac{\ell n}{k})=o(\ell\log k)$. In fact, by $U_{\mathcal{J},\mathcal{J}^{\prime},1}\sim\text{Bin}(M_{\mathcal{J}1},P_{2})$ and $M_{\mathcal{J}1}=\frac{(1-\zeta)Cn\nu e^{-\nu}\ell}{k}\leq\frac{C_{0}n\ell}{k}$ (recall that we are now considering $C\leq C_{0}=O(1)$ and $\nu e^{-\nu}\leq 1$ holds trivially), for fixed $\mathcal{J},\mathcal{J}^{\prime}$ we have
$\mathbb{P}\left(U_{\mathcal{J},\mathcal{J}^{\prime},1}\geq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}\right)$ (111)
$\leq\mathbb{P}\left(\text{Bin}\Big{(}\frac{C_{0}n\ell}{k},\frac{\nu\ell(1+o(1))}{k}\Big{)}\geq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}\right)$ (112)
$\leq\exp\left(-\frac{\ell n}{k}\Big{(}\Big{[}\log\frac{k}{\ell}\Big{]}^{-1/2}\Big{[}\log\frac{k}{C_{0}\nu\ell(\log\frac{k}{\ell})^{1/2}}+1\Big{]}-\frac{C_{0}\nu\ell}{k}\Big{)}\right)$ (113)
$=\exp\Big{(}-\Omega\Big{(}\sqrt{\log\frac{k}{\ell}}\Big{)}\cdot\ell\log k\Big{)},$ (114)
@@
where we use Chernoff bound (see (24) in Lemma 1) in (113), and (114) holds since $\frac{\ell n}{k}=\Theta(\ell\log k)$. By a union bound over at most $\binom{k}{\ell}\times\binom{p}{\ell}$ choices of $(\mathcal{J},\mathcal{J}^{\prime})$, we obtain that
@@
$\mathbb{P}\left(\max_{\mathcal{J},\mathcal{J}^{\prime}}U_{\mathcal{J},\mathcal{J}^{\prime},1}\geq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}\right)\leq\exp\Big{(}\ell\log\frac{ek}{\ell}+\ell\log\frac{ep}{\ell}-\Omega\Big{(}\sqrt{\log\frac{k}{\ell}}\Big{)}\cdot\ell\log k\Big{)}\leq k^{-10\ell},$ (115)
@@
where the last inequality holds for large enough $k$ because $\frac{k}{\ell}\to\infty$ and $k=\Theta(p^{\theta})$ for some $\theta\in(0,1)$. Combining with $\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}=o(\frac{\ell n}{k})$, with probability at least $1-k^{-10\ell}$ we have
$\max_{\mathcal{J},\mathcal{J}^{\prime}}\leavevmode\nobreak\ U_{\mathcal{J},\mathcal{J}^{\prime},1}=o\Big{(}\frac{\ell n}{k}\Big{)},$ (116)
under which the sufficient condition for (C2) given in (110) can be further simplified to
@@
$\max_{\begin{subarray}{c}\mathcal{J}^{\prime}\subset[p]\setminus s\\ |\mathcal{J}^{\prime}|=\ell\end{subarray}}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta-o(1))\frac{Cn\nu e^{-\nu}\ell}{k}.$ (117)
@@
Decomposing the event (117): With respect to the randomness of $\mathbf{X}_{s}$ and $\mathbf{Y}$, we have already shown that (59) holds with $1-o(1)$ probability, and thus we can proceed on the events
$\mathscr{A}_{1}=\big{\{}N_{01}=\rho ne^{-\nu}(1+o(1)),\leavevmode\nobreak\ N_{00}=(1-\rho)ne^{-\nu}(1+o(1))\big{\}}$ (118)
@@
by a simple union bound. In the following, for a generic event $\mathscr{E}$, an upper bound on $\mathbb{P}(\mathscr{E})$ should be more precisely understood as a bound on $\mathbb{P}(\mathscr{E}\cap\mathscr{A}_{1})$, but to avoid cumbersome notation we will only make this explicit in the final stage (see (149) below).
@@
$\widetilde{G}_{\mathcal{J}^{\prime}}=\widetilde{G}_{\mathcal{J}^{\prime},1}-\widetilde{G}_{\mathcal{J}^{\prime},2}\sim\text{Bin}(\rho ne^{-\nu}(1+o(1)),P_{2})-\text{Bin}((1-\rho)ne^{-\nu}(1+o(1)),P_{2}),$ (119)
where $P_{2}$ is the probability defined in (105). We can now introduce
$d_{0}:=\frac{C(1-2\zeta)+1-\rho}{\rho}$ (120)

### 23 Oct 2024 v2

@@ Appendix D $-\ell$ Achievability Analysis for Bernoulli Designs
In this part, we identify the more explicit condition for (C1) under the given $\ell$ and a specific feasible $(C,\zeta)$ with $C\leq C_{0}$. We seek to bound $k_{\ell,C,\zeta}$. Our strategy is to first calculate $\mathbbm{E}k_{\ell,C,\zeta}$ and then deduce the high-probability behavior of $k_{\ell,C,\zeta}$ via Markov’s inequality. Since $k_{\ell,C,\zeta}=\sum_{\mathcal{J}\subset s,|\mathcal{J}|=\ell}\mathbbm{1}(M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k},\leavevmode\nobreak\ M_{\mathcal{J}0}=\zeta M_{\mathcal{J}})$, we have
$\mathbbm{E}k_{\ell,C,\zeta}$ $=\binom{k}{\ell}\mathbb{P}\left(\text{for fixed }\mathcal{J}\subset s\text{ with }|\mathcal{J}|=\ell,\leavevmode\nobreak\ M_{\mathcal{J}}=\frac{Cn\nu e^{-\nu}\ell}{k},\leavevmode\nobreak\ M_{\mathcal{J}0}=\zeta M_{\mathcal{J}}\right)$ (D.9)
$=\binom{k}{\ell}\mathbb{P}\left(\text{Bin}\big{(}n,P_{1}\big{)}=\frac{Cn\nu e^{-\nu}\ell}{k}\right)\mathbb{P}\left(\text{Bin}\Big{(}\frac{Cn\nu e^{-\nu}\ell}{k},\rho\Big{)}=\frac{\zeta\cdot Cn\nu e^{-\nu}\ell}{k}\right)$ (D.10)
$\leq\binom{k}{\ell}\exp\left(-n\cdot D\Big{(}\frac{C\nu e^{-\nu}\ell}{k}\big{\|}\frac{\nu e^{-\nu}\ell(1+o(1))}{k}\Big{)}-\frac{Cn\nu e^{-\nu}\ell}{k}D(\zeta\|\rho)\right)$ (D.11)
$=\binom{k}{\ell}\exp\Big{(}-\frac{n\nu e^{-\nu}\ell}{k}\Big{[}f_{1}(C,\zeta,\rho)+o(1)\Big{]}\Big{)},$ (D.12)
where in (D.11) we use the Chernoff bound (see (A.1) in Lemma A.1), and in (D.12) we define
$f_{1}(C,\zeta,\rho):=C\log C-C+C\cdot D(\zeta\|\rho)+1,$ (D.13)
and then use $D(\frac{C\nu e^{-\nu}\ell}{k}\|\frac{\nu e^{-\nu}\ell(1+o(1))}{k})=\frac{\nu e^{-\nu}\ell}{k}(C\log C-C+1+o(1))$. We note that $f_{1}(C,\zeta,\rho)$ coincides with (C.20) in the proof of the converse bound for Bernoulli designs.
@@
$\mathbb{P}\Big{(}k_{\ell,C,\zeta}\geq(\ell\log k)^{5}\mathbbm{E}[k_{\ell,C,\zeta}]\Big{)}\leq(\ell\log k)^{-5}.$ (D.14)
Using the bound on $\mathbbm{E}k_{\ell,C,\zeta}$ from (D.12), we know that with probability at least $1-(\ell\log k)^{-5}$, it holds that
$k_{\ell,C,\zeta}<(\ell\log k)^{5}\mathbbm{E}k_{\ell,C,\zeta}\leq$ $(\ell\log k)^{5}\binom{k}{\ell}\exp\Big{(}-\frac{n\nu e^{-\nu}\ell}{k}\Big{[}f_{1}(C,\zeta,\rho)+o(1)\Big{]}\Big{)}$ (D.15)
$\leq$ $\exp\left((1+o(1))\ell\log\frac{k}{\ell}-\frac{n\nu e^{-\nu}\ell}{k}\Big{[}f_{1}(C,\zeta,\rho)+o(1)\Big{]}\right).$ (D.16)
@@
$\mathbb{P}\Big{(}k_{\ell,C,\zeta}\geq k^{5}\mathbbm{E}[k_{\ell,C,\zeta}]\Big{)}\leq k^{-5}.$ (D.17)
Combining with (D.12), we know that with probability at least $1-k^{-5}$, it holds that
$k_{\ell,C,\zeta}<k^{5}\mathbbm{E}k_{\ell,C,\zeta}\leq$ $k^{5}\binom{k}{\ell}\exp\Big{(}-\frac{n\nu e^{-\nu}\ell}{k}\Big{[}f_{1}(C,\zeta,\rho)+o(1)\Big{]}\Big{)}$ (D.18)
$\leq$ $\exp\Big{(}(1+o(1))\ell\log\frac{k}{\ell}-\frac{n\nu e^{-\nu}\ell}{k}\Big{[}f_{1}(C,\zeta,\rho)+o(1)\Big{]}\Big{)}.$ (D.19)
@@
Combining the above two cases, we obtain the following implication for given $(\ell,C,\zeta)$ for any $\eta_{1}>0$ and sufficiently large $k$:
$\frac{n\nu e^{-\nu}\ell}{k}\big{[}f_{1}(C,\zeta,\rho)+o(1)\big{]}\geq(1+\eta_{1})\ell\log\frac{k}{\ell}$ (D.20)
@@
$\Longrightarrow\leavevmode\nobreak\ \mathbb{P}\big{(}k_{\ell,C,\zeta}\geq 1\big{)}\leq\hat{P}_{\ell},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{where }\hat{P}_{\ell}=\begin{cases}(\ell\log k)^{-5},\leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }1\leq\ell\leq\log k\\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ k^{-5},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \text{if }\log k<\ell\leq\frac{k}{\log k}.\end{cases}$ (D.21)
@@
Therefore, if (D.20) holds, then (C1) holds for some given $(\ell,C,\zeta)$ with probability at most $\hat{P}_{\ell}$.
Condition for (C2) for a given $(\ell,C,\zeta)$
@@
We again consider a given $\ell$ and a specific feasible pair $(C,\zeta)$ with $C\leq C_{0}$, and now switch to analyzing (C2). For $\mathcal{J}^{\prime}\subset[p]$ with $|\mathcal{J}^{\prime}|=\ell$, the probability of a test containing some item from $\mathcal{J}^{\prime}$ is:
$P_{2}:=1-\big{(}1-\frac{\nu}{k}\big{)}^{\ell}=\frac{\nu\ell(1+o(1))}{k}.$ (D.22)
@@
Given $(\ell,C,\zeta)$, (C2) states that (D.1) holds for some $\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}$ and some $\mathcal{J}^{\prime}\subset[p]\setminus s$ with $|\mathcal{J}^{\prime}|=\ell$. In (D.1), recall that $G_{\mathcal{J},\mathcal{J}^{\prime},1}$ (resp., $G_{\mathcal{J},\mathcal{J}^{\prime},2}$) is the number of tests in $\mathcal{N}_{01}\cup\mathcal{M}_{\mathcal{J}1}$ (resp., $\mathcal{N}_{00}\cup\mathcal{M}_{\mathcal{J}0}$) that contain some item from $\mathcal{J}^{\prime}$, and it will be convenient to further decompose these into
$G_{\mathcal{J},\mathcal{J}^{\prime},1}=\widetilde{G}_{\mathcal{J}^{\prime},1}+U_{\mathcal{J},\mathcal{J}^{\prime},1},\leavevmode\nobreak\ G_{\mathcal{J},\mathcal{J}^{\prime},2}=\widetilde{G}_{\mathcal{J}^{\prime},2}+U_{\mathcal{J},\mathcal{J}^{\prime},2},$ (D.23)
@@
where we define $\widetilde{G}_{\mathcal{J}^{\prime},1}$ (resp., $\widetilde{G}_{\mathcal{J}^{\prime},2}$) as the number of tests in $\mathcal{N}_{01}$ (resp., $\mathcal{N}_{00}$) that contain some item from $\mathcal{J}^{\prime}$ (thus $\widetilde{G}_{\mathcal{J}^{\prime},1}$ and $\widetilde{G}_{\mathcal{J}^{\prime},2}$ have no dependence on $\mathcal{J}$, as also reflected by our notation), and $U_{\mathcal{J},\mathcal{J}^{\prime},1}$ (resp., $U_{\mathcal{J},\mathcal{J}^{\prime},2}$) as the number of tests in $\mathcal{M}_{\mathcal{J}1}$ (resp., $\mathcal{M}_{\mathcal{J}0}$) that contain some item from $\mathcal{J}^{\prime}$. Given $(\mathbf{X}_{s},\mathbf{Y})$ and for specific $\mathcal{J}$ and $\mathcal{J}^{\prime}$, it follows that
$\widetilde{G}_{\mathcal{J}^{\prime},1}\sim\text{Bin}(N_{01},P_{2}),\leavevmode\nobreak\ \widetilde{G}_{\mathcal{J}^{\prime},2}\sim\text{Bin}(N_{00},P_{2})$ (D.24)
$U_{\mathcal{J},\mathcal{J}^{\prime},1}\sim\text{Bin}(M_{\mathcal{J}1},P_{2}),\leavevmode\nobreak\ U_{\mathcal{J},\mathcal{J}^{\prime},2}\sim\text{Bin}(M_{\mathcal{J}0},P_{2})$ (D.25)
@@
are independent, as the tests comprising $\mathcal{N}_{01},\mathcal{N}_{00},\mathcal{M}_{\mathcal{J}0},\mathcal{M}_{\mathcal{J}1}$ are disjoint. Under these definitions, (D.1) can be written as
$\underbrace{\widetilde{G}_{\mathcal{J}^{\prime},1}-\widetilde{G}_{\mathcal{J}^{\prime},2}}_{:=\widetilde{G}_{\mathcal{J}^{\prime}}}+U_{\mathcal{J},\mathcal{J}^{\prime},1}-U_{\mathcal{J},\mathcal{J}^{\prime},2}\geq(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k},$ (D.26)
@@
$\max_{\begin{subarray}{c}\mathcal{J}^{\prime}\subset[p]\setminus s\\ |\mathcal{J}^{\prime}|=\ell\end{subarray}}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta)\frac{Cn\nu e^{-\nu}\ell}{k}-\max_{\begin{subarray}{c}\mathcal{J}\in\mathcal{K}_{\ell,C,\zeta}\\ \mathcal{J}^{\prime}\subset[p]\setminus s,|\mathcal{J}^{\prime}|=\ell\end{subarray}}\leavevmode\nobreak\ U_{\mathcal{J},\mathcal{J}^{\prime},1}.$ (D.27)
Thus, to bound the probability of (C2), it suffices to bound the probability of its necessary condition given in (D.27).
@@
The effect of $U_{\mathcal{J},\mathcal{J}^{\prime},1}$: For convenience, we write the last term in (D.27) using the shorthand $\max_{\mathcal{J},\mathcal{J}^{\prime}}U_{\mathcal{J},\mathcal{J}^{\prime},1}$ with the constraints left implicit. We first show that the term $\max_{\mathcal{J},\mathcal{J}^{\prime}}U_{\mathcal{J},\mathcal{J}^{\prime},1}$ in the right-hand side of (D.27) only has a minimal impact, in the sense that scales as $o(\frac{\ell n}{k})=o(\ell\log k)$. In fact, by $U_{\mathcal{J},\mathcal{J}^{\prime},1}\sim\text{Bin}(M_{\mathcal{J}1},P_{2})$ and $M_{\mathcal{J}1}=\frac{(1-\zeta)Cn\nu e^{-\nu}\ell}{k}\leq\frac{C_{0}n\ell}{k}$ (recall that we are now considering $C\leq C_{0}=O(1)$ and $\nu e^{-\nu}\leq 1$ holds trivially), for fixed $\mathcal{J},\mathcal{J}^{\prime}$ we have
$\mathbb{P}\left(U_{\mathcal{J},\mathcal{J}^{\prime},1}\geq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}\right)$
$\leq\mathbb{P}\left(\text{Bin}\Big{(}\frac{C_{0}n\ell}{k},\frac{\nu\ell(1+o(1))}{k}\Big{)}\geq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}\right)$ (D.28)
$\leq\exp\left(-\frac{\ell n}{k}\Big{(}\Big{[}\log\frac{k}{\ell}\Big{]}^{-1/2}\Big{[}\log\frac{k}{C_{0}\nu\ell(\log\frac{k}{\ell})^{1/2}}-1\Big{]}+\frac{C_{0}\nu\ell}{k}\Big{)}\right)$ (D.29)
$=\exp\Big{(}-\Omega\Big{(}\sqrt{\log\frac{k}{\ell}}\Big{)}\cdot\ell\log k\Big{)},$ (D.30)
@@
where in (D.29) we use the Chernoff bound (see (A.2) in Lemma A.1), and (D.30) holds since $\frac{\ell n}{k}=\Theta(\ell\log k)$. By a union bound over at most $\binom{k}{\ell}\times\binom{p}{\ell}$ choices of $(\mathcal{J},\mathcal{J}^{\prime})$, we obtain that
@@
$\mathbb{P}\left(\max_{\mathcal{J},\mathcal{J}^{\prime}}U_{\mathcal{J},\mathcal{J}^{\prime},1}\geq\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}\right)\leq\exp\Big{(}\ell\log\frac{ek}{\ell}+\ell\log\frac{ep}{\ell}-\Omega\Big{(}\sqrt{\log\frac{k}{\ell}}\Big{)}\cdot\ell\log k\Big{)}\leq k^{-10\ell},$ (D.31)
@@
where the last inequality holds for large enough $k$ because $\frac{k}{\ell}\to\infty$ and $k=\Theta(p^{\theta})$ for some $\theta\in(0,1)$. Combining with $\frac{\ell n}{k(\log\frac{k}{\ell})^{1/2}}=o(\frac{\ell n}{k})$, with probability at least $1-k^{-10\ell}$, we have
$\max_{\mathcal{J},\mathcal{J}^{\prime}}\leavevmode\nobreak\ U_{\mathcal{J},\mathcal{J}^{\prime},1}=o\Big{(}\frac{\ell n}{k}\Big{)},$ (D.32)
under which the sufficient condition for (C2) given in (D.27) can be further simplified to
@@
$\max_{\begin{subarray}{c}\mathcal{J}^{\prime}\subset[p]\setminus s\\ |\mathcal{J}^{\prime}|=\ell\end{subarray}}\widetilde{G}_{\mathcal{J}^{\prime}}\geq(1-2\zeta-o(1))\frac{Cn\nu e^{-\nu}\ell}{k}.$ (D.33)
@@
Decomposing the event (D.33): With respect to the randomness of $\mathbf{X}_{s}$ and $\mathbf{Y}$, we have already shown that (C.23) holds with $1-o(1)$ probability, and thus we can proceed on the event
$\mathscr{A}_{1}=\big{\{}N_{01}=\rho ne^{-\nu}(1+o(1)),\leavevmode\nobreak\ N_{00}=(1-\rho)ne^{-\nu}(1+o(1))\big{\}}$ (D.34)
@@
by a simple union bound. In the following, for a generic event $\mathscr{E}$, an upper bound on $\mathbb{P}(\mathscr{E})$ should be more precisely understood as a bound on $\mathbb{P}(\mathscr{E}\cap\mathscr{A}_{1})$, but to avoid cumbersome notation we will only make this explicit in the final stage (see (D.64) below).
@@
$\widetilde{G}_{\mathcal{J}^{\prime}}=\widetilde{G}_{\mathcal{J}^{\prime},1}-\widetilde{G}_{\mathcal{J}^{\prime},2}\sim\text{Bin}(\rho ne^{-\nu}(1+o(1)),P_{2})-\text{Bin}((1-\rho)ne^{-\nu}(1+o(1)),P_{2}),$ (D.35)
where $P_{2}$ is the probability defined in (D.22). We can now introduce
$d_{0}:=\frac{C(1-2\zeta)+1-\rho}{\rho}$ (D.36)